Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability of once in four

Expert replies
by rajataga » Sun Jan 11, 2009 8:07 am
A man can hit a target once in 4 shots. If he fires 4 shots in succession, what is the probability that he will hit his target?

A. 1

B. 255/256

C. 175/256

D. 1/4

E. 1/2
Join the discussion
Source: — Problem Solving |

Re: Probability of once in four

by Stuart@KaplanGMAT » Sun Jan 11, 2009 8:13 am
rajataga wrote:A man can hit a target once in 4 shots. If he fires 4 shots in succession, what is the probability that he will hit his target?

A. 1

B. 255/256

C. 175/256

D. 1/4

E. 1/2
There are 4 different scenarios in which he will "hit his target".

S1: hit once, miss 3 times
S2: hit twice, miss twice
S3: hit 3 times, miss once
S4: hit all 4 times

One way we could answer this question is to calculate the probability of each scenario and add them together.

However, a much quicker way to answer the question is to recognize that there's only 1 scenario we do NOT want: miss all 4 times. We can use our trusty "one minus" probability formula:

Prob(want) = 1 - Prob(don't want)

So, the probability of missing all 4 is:

3/4 * 3/4 * 3/4 * 3/4 = 81/256

Therefore, the probability of hitting at least one target is:

1 - 81/256 = 175/256... choose (C).
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by rajataga » Sun Jan 11, 2009 8:21 am
Thanks Stuart, you are correct.

However, what is the reason that in this scenario, you cannot think of them as 4 exclusive events, and say that if he takes 4 shots at the target, he will hit it atleast once?
Join the discussion

by rajataga » Sun Jan 11, 2009 8:35 am
i guess i know the answer mathematically....

each of these events are exclusive, and the outcome of one doesn't affect the outcome of another.....However, i am not yet convinced about this....
Join the discussion

by Stuart@KaplanGMAT » Sun Jan 11, 2009 9:28 am
rajataga wrote:i guess i know the answer mathematically....

each of these events are exclusive, and the outcome of one doesn't affect the outcome of another.....However, i am not yet convinced about this....
Right.. the events are independent, which means the outcome of one doesn't affect the outcome of another.

Think about rolling a 6-sided die. There's a 1 in 6 chance of rolling each number. Does that mean that if you roll the die 6 times, you'll always get a 1, 2, 3, 4, 5 and 6? Of course not!

Of course, part of the problem is the wording of the question you posted, which is far too ambigious for an actual GMAT question. On the actual test, the question would likely have read something along the lines of:

"The probability of a man hitting a target with any shot he takes is 1/4. If the man fires 4 shots at the target, what's the chance that he hits it at least once?"

What's the source of the question?
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by rajataga » Sun Jan 11, 2009 10:47 am
Came across it on 4gmat.com....

Actually very true, the question is kinda worded wrong if u think about the first part...
Join the discussion

by Zizou4 » Thu Jan 22, 2009 2:04 am
After considering the scenarios 1 by 1 and adding them up, I still dont get it right. Could someone plse check my formula?

P(TMMM) + P(TTMM) + P(TTTM) + P(TTTT)

WHERE:
P(TMMM)= (1/4) * (3/4)^1/3 * 4(ways of hitting target once)
P(TTMM)= (1/4) ^1/2 * (3/4)^1/2 * 6
P(TTTM)= (1/4) ^1/3 * (3/4) * 4
P(TTTT)= (1/4) ^1/4 * 1

I am sure the problem lies in my "ways of hitting target" calculation.

Thanx in advance.
You do not have the required permissions to view the files attached to this post.
Join the discussion